OpenStudy (anonymous):

If f(x)=3x and g(x)=2x-6 find (g o g)(x)

OpenStudy (anonymous):

g(g(x)) ?? Put x = 2x - 6 in g(x)..

OpenStudy (unklerhaukus):

\[(g\circ g)(x)=g(g(x))=g(2x-6)=dots\]

OpenStudy (anonymous):

for the "dots' replace the "\(x\)" in \(2x-6\) by \(2x-6\)

OpenStudy (anonymous):

Replace the x with what ?

OpenStudy (anonymous):

or if you prefer, replace the \(\diamondsuit\) in \(2\diamondsuit -6\) by \(2x-6\)

OpenStudy (anonymous):

let me write it out for you \[g(g(x))=g(2x-6)=2(2x-6)-6\] then some algebra to clean it up

OpenStudy (anonymous):

celeste wants to have her hair cut and permedand also go to lunch .she knows she will need $59 . the perm cost twice as much as her hair cut and she needs $5 for lunch . how much does the perm cost ?

OpenStudy (anonymous):

don't let all the "\(x\)'s confuse you the function is not "two ex minus six" the function says "take the input, multiply it by 2, then subtract 6"

OpenStudy (anonymous):

so if i have \(g(5)\) it is \(2\times 5-6\) and if i have \(g(\heartsuit)\) it is \(2(\heartsuit)-6\) and \[g(\heartsuit +\clubsuit)=2(\heartsuit +\clubsuit)-6\] and so \[g(2x-6)=2(2x-6)-6\]

OpenStudy (anonymous):

no my problem is :celeste wants to have her hair cut and permedand also go to lunch .she knows she will need $59 . the perm cost twice as much as her hair cut and she needs $5 for lunch . how much does the perm cost ?

OpenStudy (anonymous):

Okay thank you. i figured it out. you were very helpful;. you deserve a medal !!!